OpenStudy (anonymous):

solve for x: 1/2+4/2x=x+4/10

OpenStudy (anonymous):

Equation editor, or brackets. Please.

OpenStudy (king):

x=6/5

OpenStudy (anonymous):

how??

OpenStudy (anonymous):



OpenStudy (king):

oh!

OpenStudy (king):

then answer is x=-4

OpenStudy (anonymous):

x=5, -4.. it will form a quadratic

OpenStudy (mimi_x3):

\[\frac{x+4}{2x} = \frac{x+4}{10} \] \[10(x+4) = 2x(x+4) \] \[10x + 40 = 2x^2 + 8x \] Then solve for x.

OpenStudy (anonymous):

I dont know that. Sorry Kyyy.